OpenStudy (abdela25):

write an equation in slope intercept form for the line that passes through the given points (5, -2), (-16, 4) MEDALS WILL BE REWARDED !!!

OpenStudy (anonymous):

y-mx+b and y2-y1/x2-x1

OpenStudy (anonymous):

m=-2/7

OpenStudy (anonymous):

y=2(x+2)/7 y=-2x/7 -4/7

OpenStudy (abdela25):

is that it or does it go more

OpenStudy (anonymous):

\[y=\frac{ -2 }{ 7 }\frac{ -4 }{ 7 }\]

OpenStudy (anonymous):

nope that's it

OpenStudy (anonymous):

the negative sign should be in line with the division sign though

OpenStudy (abdela25):

ok thank you very much :D greatly appreciated and medal was awarded

OpenStudy (anonymous):

Thanks!

OpenStudy (abdela25):

ok can you help me with another problem

OpenStudy (anonymous):

I'll try

OpenStudy (abdela25):

ok here it is write an equation in slope-intercept form for the line that passes through the given point AND is parallel to the given line. (-3, 5); y= -1/2x+4 **theres also another set of points but i will give them to you after this question**

OpenStudy (anonymous):

The parallel line is y=-1/2x+7/2

OpenStudy (abdela25):

could you explain a little bit im kind of lost in these type of questions

OpenStudy (anonymous):

Start with the slope intercept formula: y=mx+b

OpenStudy (abdela25):

okay and then just add the numbers to replace the variables?

OpenStudy (anonymous):

so m=-1/2

OpenStudy (anonymous):

you need to find b then plug in the variables

OpenStudy (abdela25):

and b=4

OpenStudy (abdela25):

right?

OpenStudy (anonymous):

b=7/2

OpenStudy (abdela25):

what? how did you get that :/ im lost again

OpenStudy (anonymous):

x=3 and y=5 from the points that were listed

OpenStudy (abdela25):

you mean x = -3 not 3

OpenStudy (anonymous):

so if m=-1/2; x=-3 and y=5 then you are going to rework the formula to solve for b. 5=(-1/2)*(-3)+b

OpenStudy (anonymous):

yes, I meant - 3 above

OpenStudy (anonymous):

b+3/2=5

OpenStudy (anonymous):

b=7/2

OpenStudy (abdela25):

oh okay i get what you did here

OpenStudy (anonymous):

now plug in the numbers to the formula and you get y=-1/2x+7/2

OpenStudy (abdela25):

yeah i get it now

OpenStudy (abdela25):

but the second set of points isnt ordered like this one

OpenStudy (anonymous):

You didn't give me a 2nd set of points. Only 1 set and a line to be parallel to

OpenStudy (abdela25):

same question different set of pairs do you have the time to help or should i post it for someone else

OpenStudy (anonymous):

I've got a little time

OpenStudy (abdela25):

ok itll be quick i think

OpenStudy (abdela25):

write an equation in slope-intercept form for the line that passes through the given point AND is parallel to the given line. (-7, 3); x=4

OpenStudy (anonymous):

Wow - ok. That one is done the same way but way different answer. Your m=infinity and there is no Y intercept. x=-7

OpenStudy (anonymous):

So if you were graphing the parallel lines the would be crossing the x axis but not the y axis. So picture a line straight up and down

OpenStudy (abdela25):

oh ok m=infinity and x=-7 and there is no y-intercept

OpenStudy (anonymous):

yes. I posted a graph of the 2 parallel lines if that helps

OpenStudy (abdela25):

so the parallel line is x=-7 correct?

OpenStudy (abdela25):

yes it helps very much thanks for the effort and time

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're welcome

OpenStudy (abdela25):

well thanks again and i know you have to go so yeah if i need help and cant find it i know who to come to :D